# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

### v4

@@ 3 Methodology
Figure 4: Overview of the three-stage training pipeline for Ivy-xDetector, including general video understanding, detection instruction tuning, and interpretability instruction tuning.
@@
As shown in Figure 3, due to limited reasoning ability, existing models struggle to detect AIGC content and generate reliable explanations. To address this, we inject Gemini-generated detection and explanatory COT into the training process, thereby improving the model’s fine-grained artifact perception and explanation quality. This stage aims to initialize Ivy-xDetector with instruction-following and explainability capabilities. The resulting model not only performs accurate AIGC detection but also generates coherent, human-understandable rationales for its classifications.
@@
Although the fine-tuned model from Stage 1 demonstrates improved artifact awareness and interpretability, its generalization ability remains limited. To further enhance the model’s capacity for producing consistent and human-understandable explanations, we adopt the Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmathpushinglimitsmathematical]. GRPO enables reinforcement learning using only binary classification samples, achieving efficient and data-light optimization. A detailed description of GRPO and the reward functions for Ivy-xDetector can be found in Appendix 9.
We construct binary real/fake pairs from multiple datasets, each containing both authentic and synthetic images and videos. As illustrated in Figure 3, Ivy-xDetector takes the visual input and extracts the text enclosed within the <conclusion></conclusion> tags, where the predicted label is either real or fake. The quantitative results of different training stages are presented in Table 7.
@@ 4.1 Experimental Details
We primarily evaluate three closed-source models and three open-source models on our Ivy-Fake. The closed-source models are GPT-4o [gpt4_achiam2023gpt] and Gemini2.5 Flash [genimi_team2023gemini]. For the open-source models, we select models of comparable size: InternVL3.5 [chen2024internvl], Qwen2.5-VL [bai2025qwen2], MiMo-VL [coreteam2025mimovltechnicalreport], LLaVA-OneVision-1.5 [LLaVA-OneVision-1.5, lillava] and MiniCPM-V-4.5 [yao2024minicpm, yu2025minicpmv45cookingefficient]. Ivy-xDetector denotes our fully RL-trained model after the GRPO optimization stage.
@@
We report standard accuracy (Acc) and macro-averaged F1 score (F1) to assess the model’s ability to distinguish real from fake instances. For the reasoning task, we measure the similarity between the model’s reasoning process and the reference annotations using the ROUGE-L score [lin2004rouge] and BertScore [bert-score, he2021deberta], which captures the longest common subsequence between predicted and reference texts, reflecting token-level overlap. Since ROUGE-L may fail to fully capture the fidelity of reasoning steps, we adopt an LLM-as-a-judge evaluation paradigm [zheng2023judging], following the FakeBench protocol [liu2024mmfakebench], which assesses model responses along four dimensions: (1) Completeness: It reflects the extent to which the response fully addresses all aspects of the user’s question. More complete responses should incorporate information aligning well with the “golden clues” or reference answers. Incomplete or partially answered responses will receive lower scores. (2) Relevance: Measure how closely the content relates to the original annotation; (3) Level of Detail: Assess whether the response includes enough examples or elaborations; (4) Explanation: Verify the accuracy and consistency of explanations for any causes mentioned. Each response is scored using GPT-4o mini [gpt4_achiam2023gpt] under a unified evaluation prompt (see Appendix 13), which instructs the model to act as an impartial judge and assign a score from 1 to 5.
Figure 5: Ivy-xDetector performance on video detection. The model provides fine-grained spatiotemporal reasoning chains for explainable analysis.
@@
Our model was trained using the AdamW [adamw_Loshchilov2017FixingWD] optimizer with a cosine learning rate schedule. During the SFT stage, the model was trained for one epoch with a batch size of 1, which ensures stable convergence on large-scale multimodal data. In the subsequent RL stage, we adopted the GRPO algorithm, training for one epoch with the same batch size and a learning rate of $1\times 10^{-5}$. Within GRPO, the group size $n$ was set to 8, the warm-up ratio to 0.01, the temperature to 0.9. Training the SFT and RL stages required approximately 10 and 50 hours, respectively, on a system equipped with 32 A100 GPUs.
@@ 4.2 Main Results
Table 4: Comparison on the Chameleon [aide_yan2025a]. Accuracy (%) of different detectors (rows) in detecting real and fake images. For each training dataset, the first row is overall accuracy, the second row is “fake/real” class accuracy.
@@
Table 5: Comparison on the Genimage [genimage_2023]. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively. Results of other methods are reported from [aide_yan2025a].
@@
CNNSpot [cnnspot_wang2020cnn] 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
F3Net [qian2020thinking] 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
DIRE [dire_wang2023dire] 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
GenDet [zhu2023gendet] 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft [aigcdetectionbenchmark_zhong2023patchcraft] 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
FRIDA [bonechi2025thisfakedetectionsource] 85.80 85.00 85.50 84.00 87.30 83.90 86.50 84.80 85.30
AIDE [aide_yan2025a] 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
DRCT [drct] 91.50 95.00 94.40 79.40 89.20 94.70 90.00 81.70 89.50
Effort [effort] 82.40 99.80 99.80 78.70 93.30 97.40 91.70 77.60 91.10
ThinkFake-7B [huang2025thinkfakereasoningmultimodallarge] 92.50 93.10 95.30 73.10 87.40 93.60 66.20 70.80 84.00
@@
Table 6: Comparisons to the GenVideo [demamba]. F1 score (F1), recall score (R) and average precision (AP) on the many-to-many generalization task. “Demamba-XCLIP-FT” is abbreviated as “Demamba”. Results of other methods are reported from [demamba].
@@ 4.3 Generalization Evaluation
We conduct evaluations across multiple public leaderboards, including the image-based GenImage 5 and the video-based GenVideo benchmark 6. Performance results of other competing detectors are taken from AIDE [aide_yan2025a] and Demamba [demamba]. Our method achieves superior performance on all these leaderboards, demonstrating superior detection accuracy and robust cross-modal consistency.
@@
Beyond the datasets aligned with our training sets, we further evaluate the generalization ability of Ivy-xDetector on the Chameleon benchmark[aide_yan2025a], which lies outside the training data distribution, as shown in Table 4. The results confirm that our approach preserves strong generalization capability across unseen generative models and diverse data domains.
@@
4.4 Case Study
As illustrated in Figure 5, Ivy-xDetector demonstrates a rich single frame / video understanding across multiple dimensions, effectively capturing subtle differences in generated content. To further highlight its advantages over other models, we provide additional comparison cases with GPT-4o, Gemini-2.5, InternVL3, and Qwen2.5-VL in Appendix 15.
@@ 5 Conclusion
We introduced Ivy-Fake, the first unified, large-scale dataset for explainable AIGC detection across both images and videos, featuring over 131K richly annotated training samples and 5,000 evaluation examples with natural-language reasoning, and proposed Ivy Explainable Detector ( Ivy-xDetector), a MLLM that jointly detects and explains synthetic content. Our model sets superior benchmarks in AIGC detection and explainability, and our publicly released resources provide a robust foundation for transparent, trustworthy multimodal analysis. We believe that such fine-grained explainability data Ivy-Fake can catalyze new research directions for the AIGC detection community, enabling deeper understanding of synthetic content and more principled model development.
@@
Broader impacts: The data and model can be helped with the detection of fake visual content, help to uncover fact. Potentially, this can also be used to train stronger generative model to synthesize more realistic visual content.
[…]
Supplementary Material
6 Author and Affiliation List
$\pi^{3}Lab$    Changjiang Jiang*, Wenhui Dong*🖂, Zhonghao Zhang
Wuhan University    Changjiang Jiang*, Fengchang Yu
Nanjing University    Chenyang Si, Caifeng Shan
Stanford University    Wei Peng
Nankai University    Xinbin Yuan
Georgia Institute of Technology    Yifei Bi
Jilin University    Ming Zhao
Zhejiang University    Zian Zhou
* Equal contribution. 🖂 Corresponding author.
7 License
For all datasets in our experiments and datasets, FLUX.1 [labs2025flux1kontextflowmatching] and Lora Flux [labs2025flux1kontextflowmatching] falls under the FLUX.1 [dev] Non-Commercial License.WildFake [wildfake_2025] is released under an open-source license, making it freely available for research and non-commercial use; Qwen-Image [wu2025qwenimagetechnicalreport] is licensed under Apache 2.0; Sora [sora_videoworldsimulators2024] is a commercial, non-open-source text-to-video model provided via ChatGPT Plus/Pro; user-generated content is owned by the user and allows (subject to usage policies) non-commercial use, but the model itself is closed-source.
Stable Diffusion [svd_blattmann2023stable], while originally open-source under Creative ML OpenRAIL-M (e.g., versions 1.X, 2.1, SDXL), has transitioned in newer versions (SD3.x, SD3.5) to the Stability AI Community License, which permits free use for individuals or entities with annual revenue under USD 1 million, but requires a paid Enterprise License for larger-scale or commercial usage.
During Ivy-Fake construction and experimentation, we accessed Gemini API and ChatGPT API to ensure annotation and evaluation’s quality and reproducibility. We explicitly state that all data generated using these commercial APIs are used solely for academic research and non-commercial purposes, fully complying with the respective API usage agreements and ethical guidelines. No Gemini- or ChatGPT-derived content is redistributed for any commercial training, deployment, or monetization purposes.
Image Authenticity Analyst Assistant (User Prompt) Is this image real or fake?
Figure 6: User Prompt Template for Image Data Distillation
Video Authenticity Analyst Assistant (User Prompt) Is this video real or fake?
Figure 7: User Prompt Template for Video Data Distillation
Image Authenticity Analyst Assistant (System Prompt) ## Role
Expert AI system for detecting image by analyzing visual anomalies across spatial plausibility.
## Analysis Dimensions
### Spatial Features: static anomaly detection
- Impractical Luminosity
 - Scene brightness measurement
 - Invisible light source detection (physical validation)
- Localized Blur
 - Focus distribution mapping (sharpness gradient)
 - Artificial depth-of-field identification (algorithmic artifacts)
- Illegible Letters
 - OCR text extraction
 - Character structural integrity (stroke continuity)
- Distorted Components
 - Anatomical/proportional accuracy (biological/object logic)
 - Physics compliance (material/gravity validation)
- Omitted Components
 - Object completeness check (edge/detail absence)
 - Partial rendering artifact detection (AI-generated traces)
- Spatial Relationships
 - Contextual object placement (scene plausibility)
 - Perspective consistency (geometric projection)
- Chromatic Irregularity
 - Color database comparison (natural distribution)
 - Unnatural hue detection (oversaturation/abrupt gradients)
- Abnormal Texture
 - Surface pattern regularity (texture repetition)
 - Material property coherence (reflectance/roughness validation)
## Reasoning Step
1. Spatial Analysis
 - Analyze static features (e.g., lighting, text, objects)
2. Conclusion: Only real or fake.
 - real: Contains verifiable capture device signatures and natural physical imperfections.
 - fake: Exhibits synthetic fingerprints including but not limited to over-regularized textures and non-physical light interactions.
The assistant first thinks about the reasoning step in the mind and then provides the user with the reason. The reasoning step and conclusion are enclosed within <think> </think> and <conclusion> </conclusion> tags, respectively, i.e., <think> reasoning step here </think> <conclusion> real or fake </conclusion>. <conclusion> content must strictly align with the user-provided authenticity label (real/fake) in both value and semantic context.
Figure 8: System Prompt Template for Image Data Distillation
Video Authenticity Analyst Assistant (System Prompt) ## Role
Expert AI system for detecting videos by analyzing visual anomalies across temporal coherence (inter-frame dynamics) and spatial plausibility (intra-frame logic).
## Analysis Dimensions
### 1. Temporal Features: Multi-frame dynamic anomaly detection - Luminance Discrepancy
 - Shadow direction consistency (cross-frame comparison)
 - Light source coordination (temporal validation)
- Awkward Facial Expression
 - Facial muscle motion continuity (expression dynamics)
 - Emotion-context alignment (temporal coherence)
- Duplicated Components
 - Repeating element pattern recognition (cross-frame tracking)
 - Natural variation analysis (sequence validation)
- Non-Spatial Relationships
 - Object interaction physics (motion trajectory validation)
 - Fusion/penetration anomalies (temporal detection)
### 2. Spatial Features: Single-frame static anomaly detection
- Impractical Luminosity
 - Scene brightness measurement (single-frame analysis)
 - Invisible light source detection (physical validation)
- Localized Blur
 - Focus distribution mapping (sharpness gradient)
 - Artificial depth-of-field identification (algorithmic artifacts)
- Illegible Letters
 - OCR text extraction (single-frame recognition)
 - Character structural integrity (stroke continuity)
- Distorted Components
 - Anatomical/proportional accuracy (biological/object logic)
 - Physics compliance (material/gravity validation)
- Omitted Components
 - Object completeness check (edge/detail absence)
 - Partial rendering artifact detection (AI-generated traces)
- Spatial Relationships
 - Contextual object placement (scene plausibility)
 - Perspective consistency (geometric projection)
- Chromatic Irregularity
 - Color database comparison (natural distribution)
 - Unnatural hue detection (oversaturation/abrupt gradients)
- Abnormal Texture
 - Surface pattern regularity (texture repetition)
 - Material property coherence (reflectance/roughness validation)
## Reasoning Step
1. Temporal Analysis
 - Track dynamic features across frames (e.g., shadows, expressions)
2. Spatial Analysis
 - Analyze static features per frame (e.g., lighting, text, objects)
3. Conclusion: Only real or fake.
 - real: Contains verifiable capture device signatures and natural physical imperfections.
 - fake: Exhibits synthetic fingerprints including but not limited to over-regularized textures and non-physical light interactions.
The assistant first thinks about the reasoning step in the mind and then provides the user with the reason. The reasoning step and conclusion are enclosed within <think> </think> and <conclusion> </conclusion> tags, respectively, i.e., <think> reasoning step here </think> <conclusion> real or fake </conclusion>. <conclusion> content must strictly align with the user-provided authenticity label (real/fake) in both value and semantic context.
Figure 9: System Prompt Template for Video Data Distillation
8 Related Work
8.1 Methods for Synthetic Content Detection
Due to growing concerns about the misuse of synthetic data [deng2024survey], research on AI-generated content (AIGC) detection has expanded rapidly in recent years. Most existing models for AI-generated images and videos formulate the task as binary classification, simply predicting whether the content is ”real” or ”fake.” Representative examples include CNN-based AIGVDet [aigvdet_bai2024ai], CNNSpot [cnnspot_wang2020cnn] and Transformer-based models such as DIRE [dire_wang2023dire] and AIDE [aide_yan2025a]. Meanwhile, several works have explored the application of multimodal large language models (MLLMs) to AIGC detection, including Synartifact [cao2024synartifact] and Bi-LORA [keita2025bi]. However, these approaches largely overlook the importance of interpretability in AIGC detection.
Some efforts attempt to introduce interpretability by leveraging spatial annotations [dong2022explaining] or frequency-domain artifact analysis [zhang2023perceptual]. Nevertheless, the resulting explanations are often difficult for humans to comprehend, as they lack clarity in natural language. This limitation is particularly evident in the video domain, where AI-generated content frequently exhibits obvious flaws, e.g., incoherent frame transitions and object inconsistency, that are easily noticed and reasoned about by humans [deng2024survey]. FakeClue [fakeclue_wen2025spot] introduces the use of vision-language models (VLMs) to provide interpretability for image-level detection, but it does not offer a unified framework that integrates both images and videos.
8.2 Datasets for Synthetic Content Detection.
Early datasets for synthetic content detection, such as CNNSpot [cnnspot_wang2020cnn], primarily collected fake images generated by GAN-based models [gan1_goodfellow2014generative, gan2_zhu2017unpaired, gan3_brock2018large]. However, with the advent of more advanced generative architectures like diffusion models [diff1_ho2020denoising, diff2_dhariwal2021diffusion, diff3_rombach2022high, diff4_hertz2022prompt, diff5_nichol2021glide] and their variants, the authenticity of generated content has significantly increased, making it more challenging for detection models to discern. This has spurred the development of newer datasets, including ArtiFact [cao2024synartifact], GenImage [genimage_2023], and WildFake [wildfake_2025]. GenImage [genimage_2023] comprises images from the 1000 ImageNet [imagenet_ILSVRC15] categories, generated by eight state-of-the-art generators such as Stable Diffusion [diff3_rombach2022high] and Midjourney. Nevertheless, these datasets predominantly focus on image-based content. More recently, datasets emphasizing interpretability have also been introduced. FakeClue [fakeclue_wen2025spot] contains a large amount of image data with explainability annotations but lacks video data. LOKI [loki_ye2024loki] offers data across 26 different categories and includes 18,000 distinct questions; however, its volume of multimodal data is relatively small and primarily suited for evaluation rather than comprehensive model training. Therefore, a critical gap exists for a unified benchmark encompassing both image and video modalities to rigorously evaluate the performance of contemporary AIGC detectors.
Figure 10: Final Reinforcement Learning Training Curves.
9 GRPO
Following DeepSeek-R1 [deepseekr1_guo2025deepseek], we adapt the Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical], an online RL algorithm designed to maximize the advantage of generated completions while constraining policy divergence from a reference model. We formalize our training process of Ivy-xDetectorusing GRPO below. Let $p$ denote a sampled prompt, and let $o_{1},o_{2},...,o_{n}$} be a group of completions generated by the current policy $\pi_{\theta}$. For each completion $G_{i}$, a reward $r_{i}$ is computed using a rule-based reward function. The group advantage for each completion is then calculated as:
$\hat{A}_{i,t}=\frac{r_{i}-mean(r)}{std(r)}$ (1)
Reward Model. For effective RL, we employ a rule-based reward that consists of accuracy and format rewards. We introduce a solid accuracy reward for AIGC Detection, which utilizes distinct functions to evaluate binary classification task. This allows for a more appropriate assessment based on the expected answer type.
10 Effect of Incorporating Human-Annotated Labels via gemini 2.5 pro on Accuracy
To assess the impact of human-annotated labels on model performance, we compare the accuracy of final conclusion predictions under two settings: (i) with labels incorporated via the gemini 2.5 pro, and (ii) without labels. The evaluation was conducted on about 1,000 examples from the test set.
Annotation Setting Accuracy (Acc)
With Label 1.000
Without Label 0.785
Table 8: Accuracy of conclusion prediction with and without incorporating labels.
As shown in Table 8, incorporating ground-truth labels results in a substantial performance gain, yielding perfect accuracy (1.000), compared to 0.785 without labels. The drastic performance gap suggests potential limitations in label-free or weakly supervised setups when applied to tasks requiring fine-grained semantic understanding.
11 Data Distribution
12 Additional Experiment
12.1 Synthetic Image Detection
We evaluated our AIGC detector on the GenImage [genimage_2023] and Chameleon [aide_yan2025a] benchmarks.
As shown in Table 5, GenImage comprises seven subsets generated by leading models, i.e., Midjourney, Stable Diffusion v1.4 & v1.5, ADM, GLIDE, Wukong, VQDM, and BigGAN. We compared against five state-of-the-art detectors, which are CNNSpot [cnnspot_wang2020cnn], F3Net [qian2020thinking], DIRE [dire_wang2023dire], GenDet [zhu2023gendet], PatchCraft [aigcdetectionbenchmark_zhong2023patchcraft], and AIDE [aide_yan2025a].
Notably, our detector surpasses previous state-of-the-art methods such as AIDE (86.88%) and PatchCraft (82.30%) by a large margin, even though it is trained under a unified multimodal paradigm rather than a generator-specific setting. The improvement is especially pronounced on challenging generators like ADM, GLIDE, and BigGAN, where conventional CNN- or patch-based detectors often fail to capture high-frequency inconsistencies or generalized texture patterns.
Notably, our model is trained with only 130K image samples—fewer than GenImage’s 2M+ training images—yet still achieves superior generalization performance.
12.2 Synthetic Video Detection
As shown in Table 6, we evaluate the performance of our detector on the GenVideo benchmark [demamba].
Specifically, Ivy-xDetector achieves an average F1 score of 95.26% and recall of 95.28%, substantially surpassing the prior state-of-the-art DeMamba (F1 = 90.20%). The improvement is particularly evident on challenging categories such as HotShot, Lavie, and Show-1, where previous methods tend to overfit to specific generative distributions or temporal artifacts. Our results indicate that the unified MLLM-based detection paradigm not only captures spatial inconsistencies but also learns transferable temporal-spatial representations, leading to more robust generalization across unseen generative models.
[…]
13 Prompts
Default System Prompt You are an AI-generated content detector. Given a single media (image or video), classify it as real or fake. Provide detailed reasoning inside the <think></think> tags, including your step-by-step thought process.
Your output must begin with <think>
n and end with </conclusion>.
Then output exactly one word in lowercase—either real or fake—wrapped in <conclusion></conclusion>.
Do not include any other words. If uncertain, choose the most likely class.
Figure 11: Default System Prompt for Training and Evaluation
Here we provide the prompts that are mainly used in this study. The default system prompt can be seen in Figure11. As illustrated by the following figures, there are five distillation prompts distillation we used in this paper that mainly can be divided into the following three parts:
Prompt Template for Image Data Distillation: Since image data consists of a single frame, it can be treated as a static instance. Therefore, AIGC detection mainly focuses on identifying spatial anomalies. Detail prompt can be found in Figure 6 and  8.
Prompt Template for Video Data Distillation: Compared to images, video inputs provide continuous multi-frame context. This allows for detection along both spatial and temporal anomaly dimensions. Dtail prompt can be found in Figure 7 and Figure 9.
GPT Assisted Evaluation Prompt: To assess the quality of model outputs, we design a GPT-based evaluator prompt that scores responses across four dimensions: Completeness, Relevance, Level of Detail, and Explanation. The evaluator receives a structured pair of GroundTruth and ModelOutput, each containing a <think> section (reasoning) and a <conclusion> (final judgment). The model must return a structured JSON object with integer scores (1–5) for each dimension. The prompt is provided in Figure 12.
GPT Assisted Evaluation Prompt ## Role
You are an impartial evaluator. Your task is to assess whether a model-generated response accurately and coherently matches a human-annotated reference answer.
Each input contains two structured components:
- ¡think¿: the reasoning or analytical explanation
- ¡conclusion¿: the final judgment (e.g., real or fake)
## Evaluation Dimensions
You should compare the ModelOutput to the GroundTruth, and assign integer scores from 1 to 5 (no decimals) for the following four dimensions:
1. Completeness
- Does the ModelOutput address all aspects covered in the GroundTruth?
- More complete responses should include all relevant information, especially key g̈olden clues.̈
- Incomplete or partially aligned answers should receive lower scores.
2. Relevance
- Does the ModelOutput discuss the same detection dimensions as in the GroundTruth?
- Temporal features include:
 - Luminance discrepancy
 - Duplicated components
 - Awkward facial expressions
 - Motion inconsistency
- Spatial features include:
 - Abnormal texture
 - Distorted or omitted components
 - Chromatic irregularity
 - Impractical luminosity
 - Localized blur, etc.
- Penalize if irrelevant aspects are introduced or relevant ones are missing.
3. Level of Detail
- Does the ModelOutput describe fine-grained visual cues in each dimension?
- High scores require specific subcomponent elaboration, not just general terms.
- Penalize vague or generic responses that lack specific observations.
4. Explanation
- Is the reasoning in <think> logically consistent with the <conclusion>?
- The explanation should provide clear, causally-linked justifications.
- Penalize if the conclusion contradicts the reasoning or lacks support.
Figure 12: GPT Assisted Evaluation Prompt
14 Training Detail
The RL training process of Ivy-xDetector is illustrated in Fig.10. Our GRPO training exhibits a stable increase in reward, eventually reaching 0.95. However, we also observe that both the clip ratio and the maximum response length surge sharply around 200–300 steps, suggesting that a portion of the training samples may be overly challenging for the model.
15 Case Study: Qualitative Comparison of Methods
According to Figures 131415, and 16, Ivy-xDetector consistently demonstrates superior performance in detecting both spatial and temporal anomalies, providing stronger generalization and robustness compared to existing baselines.
Figure 13: Image example 1 where Ivy-xDetector successfully detects subtle spatial anomalies missed by baselines.
Figure 14: Image example 2 illustrating improved robustness of Ivy-xDetector against visually deceptive artifacts.
Figure 15: Video example 1 showcasing Ivy-xDetector’s superior ability to detect subtle cross-frame temporal artifacts.
Figure 16: Video example 2 showcasing Ivy-xDetector’s superior ability to detect subtle cross-frame temporal artifacts.

### v5

@@ 3.5 Method
Figure 4: Overview of the three-stage training pipeline for Ivy-xDetector.
@@
Due to limited reasoning ability, existing models struggle to detect AIGC content and generate reliable explanations. To address this, we inject Gemini-generated detection and explanatory COT into the training process, thereby improving the model’s fine-grained artifact perception and explanation quality. This stage aims to initialize Ivy-xDetector with instruction following and explainability capabilities. The resulting model not only performs accurate AIGC detection but also generates coherent, human-understandable rationales for its classifications.
@@
Although the fine-tuned model from Stage 1 demonstrates improved artifact awareness and interpretability, its generalization ability remains limited. To further enhance the model’s capacity for producing consistent and human-understandable explanations, we adopt the Group Relative Policy Optimization (GRPO) algorithm (shao2024deepseekmathpushinglimitsmathematical). GRPO enables reinforcement learning (zhoubao-earlvr; zhao2026know) using only binary classification samples, achieving efficient and data-light optimization.
We construct binary real/fake pairs from multiple datasets, each containing both authentic and synthetic images and videos. Ivy-xDetector takes the visual input and extracts the text enclosed within the <conclusion></conclusion> tags, where the predicted label is either real or fake. The quantitative results of different training stages are presented in Table 7.
3.6 GRPO
Following DeepSeek-R1 (deepseekr1_guo2025deepseek) and TextShield-R1 qu2026textshield, we adapt the Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical), an online RL algorithm designed to maximize the advantage of generated completions while constraining policy divergence from a reference model. We formalize our training process of Ivy-xDetector using GRPO below. Let $p$ denote a sampled prompt, and let $o_{1},o_{2},...,o_{n}$} be a group of completions generated by the current policy $\pi_{\theta}$. For each completion $G_{i}$, a reward $r_{i}$ is computed using a rule-based reward function. The group advantage for each completion is then calculated as:
$\hat{A}_{i,t}=\frac{r_{i}-mean(r)}{std(r)}$ (1)
3.7 Reward Model
For effective RL, we employ a rule-based reward that consists of accuracy and format rewards. We introduce a solid accuracy reward for AIGC Detection, which utilizes distinct functions to evaluate binary classification task. This allows for a more appropriate assessment based on the expected answer type.
@@ 4.1 Experimental Details
We primarily evaluate three closed-source models and three open-source models on our Ivy-Fake. The closed-source models are GPT-4o (gpt4_achiam2023gpt) and Gemini2.5 Flash (genimi_team2023gemini). For the open-source models, we select models of comparable size: InternVL3.5 (chen2024internvl), Qwen2.5-VL (bai2025qwen2), MiMo-VL (coreteam2025mimovltechnicalreport), LLaVA-OneVision-1.5 (LLaVA-OneVision-1.5; lillava) and MiniCPM-V-4.5 (yao2024minicpm; yu2025minicpmv45cookingefficient). Ivy-xDetector denotes our fully RL-trained model after the GRPO optimization stage.
@@
We report standard accuracy (Acc) and macro-averaged F1 score (F1) to assess the model’s ability to distinguish real from fake instances. For the reasoning task, we measure the similarity between the model’s reasoning process and the reference annotations using the ROUGE-L score (lin2004rouge) and BertScore (bert-score; he2021deberta), which captures the longest common subsequence between predicted and reference texts, reflecting token-level overlap. Since ROUGE-L may fail to fully capture the fidelity of reasoning steps, we adopt an LLM-as-a-judge evaluation paradigm (zheng2023judging), following the FakeBench protocol (liu2024mmfakebench), which assesses model responses along four dimensions: (1) Completeness: It reflects the extent to which the response fully addresses all aspects of the user’s question. More complete responses should incorporate information aligning well with the “golden clues” or reference answers. Incomplete or partially answered responses will receive lower scores. (2) Relevance: Measure how closely the content relates to the original annotation; (3) Level of Detail: Assess whether the response includes enough examples or elaborations; (4) Explanation: Verify the accuracy and consistency of explanations for any causes mentioned. Each response is scored using GPT-4o mini (gpt4_achiam2023gpt), which instructs the model to act as an impartial judge and assign a score from 1 to 5.
Figure 5: Ivy-xDetector performance on video detection.
@@
Our model was trained using the AdamW (adamw_Loshchilov2017FixingWD) optimizer with a cosine learning rate schedule. During the SFT stage, the model was trained for one epoch with a batch size of 1, which ensures stable convergence on large-scale multimodal data. In the subsequent RL stage, we adopted the GRPO algorithm, training for one epoch with the same batch size and a learning rate of $1\times 10^{-5}$. Within GRPO, the group size $n$ was set to 8, the warm-up ratio to 0.01, the temperature to 0.9. Training the SFT and RL stages required approximately 10 and 50 hours, respectively, on a system equipped with 32 A100 GPUs.
@@ 4.2 Main Results
Table 4: Comparison on the Chameleon (aide_yan2025a). Accuracy (%) of different detectors (rows) in detecting real and fake images. For each training dataset, the first row is overall accuracy, the second row is “fake/real” class accuracy.
@@
Table 5: Comparison on the Genimage (genimage_2023). Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). The best result and the second-best result are marked in bold and underline, respectively. Results of other methods are reported from (aide_yan2025a).
@@
CNNSpot (cnnspot_wang2020cnn) 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
F3Net (qian2020thinking) 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
DIRE (dire_wang2023dire) 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
GenDet (zhu2023gendet) 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft (aigcdetectionbenchmark_zhong2023patchcraft) 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
FRIDA (bonechi2025thisfakedetectionsource) 85.80 85.00 85.50 84.00 87.30 83.90 86.50 84.80 85.30
AIDE (aide_yan2025a) 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
DRCT (drct) 91.50 95.00 94.40 79.40 89.20 94.70 90.00 81.70 89.50
Effort (effort) 82.40 99.80 99.80 78.70 93.30 97.40 91.70 77.60 91.10
ThinkFake-7B (huang2025thinkfakereasoningmultimodallarge) 92.50 93.10 95.30 73.10 87.40 93.60 66.20 70.80 84.00
@@
Table 6: Comparisons to the GenVideo (demamba). F1 score (F1), recall score (R) and average precision (AP) on the many-to-many generalization task. “Demamba-XCLIP-FT” is abbreviated as “Demamba”. Results of other methods are reported from (demamba).
@@ 4.3 Generalization Evaluation
We conduct evaluations across multiple public leaderboards, including the image-based GenImage 5 and the video-based GenVideo benchmark 6. Performance results of other competing detectors are taken from AIDE (aide_yan2025a) and Demamba (demamba). Our method achieves superior performance on all these leaderboards, demonstrating superior detection accuracy and robust cross-modal consistency.
@@
Beyond the datasets aligned with our training sets, we further evaluate the generalization ability of Ivy-xDetector on the Chameleon benchmarkaide_yan2025a, which lies outside the training data distribution, as shown in Table 4. The results confirm that our approach preserves strong generalization capability across unseen generative models and diverse data domains.
@@ 5 Conclusion
We introduced Ivy-Fake, the first unified, large-scale dataset for explainable AIGC detection across both images and videos, featuring over 106K richly annotated training samples and 5,000 evaluation examples with natural-language reasoning, and proposed Ivy Explainable Detector (Ivy-xDetector), a MLLM that jointly detects and explains synthetic content. Our model sets superior benchmarks in AIGC detection and explainability, and our publicly released resources provide a robust foundation for transparent, trustworthy multimodal analysis. We believe that such fine-grained explainability data Ivy-Fake can catalyze new research directions for the AIGC detection community, enabling deeper understanding of synthetic content and more principled model development.
@@
Broader impacts: The interpretability evaluation framework we propose is currently applied only to the field of Fake Image Detection (du2025forensichub), which is strictly a binary classification task. In future work, our multi-dimensional evaluation approach can be extended to assess the interpretability of other tasks, such as chain-of-thought evaluation for Table Question Answering (jiang2025tabdsr), question answering in medical scenarios (zhao2026spinebench; xin2024artificial; xin2026dualcpt), GUI Agents (yuanse), Document Generation ying2024dig; ying2024dig; ying2024fine and video grounding interpretability (lin2025audio).